In each heart beat, a heart pumps 80 ml of blood at an average pressure of 100 mm of Hg. What will the power output of the heart? (Assume 60 heart beat per minute)

Grade:12th pass

1 Answers

Arun
5 years ago
The work required per beat is
W = V * (Delta P)where V is the volume pumped and delta p is the pressure increase provided.
W = 80 * 10^-6 * 100 * 1.01 * 10^5 /760
W = 1.06 Joules
Time between each neat = 60/60 = 1 sec
Hence power P = W/t = 1.06 watt